# Saleh Ashkboos ETH Zurich

### v1

@@ 5 Experimental Validation
We implement QuaRot using Hugging Face (Wolf et al., 2019) on top of the PyTorch framework (Paszke et al., 2019). To quantize the inputs, we use per-token symmetric quantization (a single scale for every row) with a constant clipping ratio of 0.9 in all our experiments. We quantize the KV caches using asymmetric quantization with a group size 128 with a constant clipping ratio of 0.95. For weight quantization, we use round-to-nearest (RTN) and GPTQ (Frantar et al., 2022) with per-column (also known as per-channel) symmetric quantization, where we extract the clipping ratio using a linear search over the squared error. We use 128 samples from WikiText-2 (Merity et al., 2016) training set with 2048 sequence length as the calibration set during GPTQ quantization. On a single NVIDIA A100 GPU, modifying Llama 2-70B with QuaRot takes 5 minutes and quantizing the model with GPTQ takes a further 2 hours.
@@ 5.1 Accuracy Results
First, we evaluate the accuracy of QuaRot on the language generation task. Table 1 shows the perplexity of Llama- 2 models on WikiText-2 when we quantize the weights using GPTQ. We compare against 4-bit SmoothQuant (Xiao et al., 2023) and OmniQuant (Shao et al., 2023). We also include the QUIK (Ashkboos et al., 2023) results when they keep all the layers (including down-projection) in 4 bits. QuaRot outperforms all previous work with at most 0.63 perplexity loss (0.47 on Llama 2-70B model). QuaRot does not need re-training steps (as in OmniQuant) nor higher precision outlier features and asymmetric quantization (as in QUIK). We also apply group-wise quantization to compare against Atom (Zhao et al., 2023) on the same number of groups for weight and activations. In this setting, QuaRot doesn’t need to keep any higher precision features and related operations (like re-ordering). QuaRot outperforms Atom with 0.1 perplexity points in the 7B model. On the 13B model, we get the same perplexity number as Atom. Our results demonstrate the effectiveness of applying the Hadamard transformation in suppressing the outliers on the generation task.
Table 1: WikiText-2 perplexity results on 4-bit quantization of Llama- 2 models with 2048 sequence length. We extract the results for SmoothQuant and OmniQuant results from (Shao et al., 2023). 128G shows the group-wise quantization with group size 128.
@@
Next, we focus on evaluating QuaRot on six important zero-shot tasks: PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), LAMBADA (OpenAI) (Radford et al., 2019), and Arc (Easy and Challenge) (Clark et al., 2018). We use the LM Evaluation Harness (Gao et al., 2021) with default parameters for our experiments. Table 2 shows the accuracy of our scheme on the above tasks as well as the average score. QuaRot preserves the accuracy with at most 4.18% average score loss (1.09% for 70B model).
Table 2: Zero-shot accuracy of Llama- 2 models with 4-bit QuaRot on PIQA (PQ), WinoGrande (WG), HellaSwag (HS), Arc-Easy (A-e), Arc-Challenge (A-c), and LAMBADA (LA).
@@ 5.2 Performance Analysis
We implement QuaRot using CUDA/12.1 on top of PyTorch and use CUTLASS for performing INT-4 matrix multiplication on TensorCore (where the results will be saved in an INT32 accumulator). We implement the attention mechanism using three routines: 1) Init: During the prefill stage, this routine initializes the cache from all the key and value vectors in the prefill. The attention output during prefill is computed directly using Flash Attention (Dao et al., 2022) since we already have access to dequantized keys and values. 2) Append: During decoding, this routine is called first to quantize the current keys and values and append them to the cache. 3) Decode: Finally, this routine is called during decoding with the current query vector. The routine computes the attention output using a quantized implementation of flash attention which can load the quantized cache and compute the final value vector. In this section, we evaluate the performance of our kernels separately. Then, we evaluate the prefill and decoding stage performances.
4-bit Linear and Attention Layers.
First, we benchmark our 4-bit linear layer which involves 4-bit matrix multiplication. For a given input of FP16, the layer optionally computes the Hadamard operation, then calls the quantization kernel to quantize and save the input in a sub-byte format. In the next step, the quantized weights and input are passed to the CUTLASS 4-bit GEMM kernel. Finally, the output is dequantized and cast back to FP16. Figure 6 shows the speedup of our 4-bit layer for different layer sizes where the layer sizes match the FFN linear layer sizes in Llama- 2 models. Our 4-bit linear layer gets 3.2x speedup relative to FP16 in Llama 2-7B model, and 4.3x on the Llama 2-70B model. These numbers are for a batch size of 1, we find that scaling is approximately linear with batch size: more results in Table 12. We include the runtime with and without Hadamard operations, as ${\mathbf{W}}_{\textrm{up}}$ and ${\mathbf{W}}_{\textrm{gate}}$ do not require Hadamard transforms, whilst ${\mathbf{W}}_{\textrm{down}}$ does. We see that the Hadamard transform adds very little overhead to the forward pass at most 7% overhead.
Figure 6: Performance of 16-bit and 4-bit linear layer for 2048 sequence lengths with and without online Hadamard transformation on a NVIDIA RTX 3090 GPU, averaged over 1000 runs. The matrix sizes correspond to the linear layer sizes in Llama- 2 FFN blocks (i.e. ${\mathbf{W}}_{\textrm{down}}$). Here the batch size is 1, but the performance ratio holds for larger batches: see Appendix A.2 for detailed results.
Prefill and Decoding Stages Performance increases.
Finally, we evaluate the improvements in memory and computation from using QuaRot at different phases of inference for the Llama 2-7B model. We divide our evaluations into two separate parts: First, we evaluate the speedup of QuaRot during the prefill (time-to-first-token) phase which is usually compute-bound. Table 4 shows the speedup of QuaRot during prefill with various batch sizes. QuaRot achieves up to 2.16x speedup on Llama 2-7B . This is less than the 3.2x achieved by the linear layers alone since the majority of attention computations is still done in FP16 or FP32 even though the smaller cache size reduces IO demand. We expect further improvement by specialized attention and more kernel optimizations. We note that QuaRot’s prefill also includes quantizing the KV cache which is needed for the decoding stage.
In addition to the speedup gained during pre-fill, we observed significant reductions in memory usage up to 3.71x during decoding with QuaRot. This improvement comes from both quantizing the KV cache as well as quantizing the model’s weights. To demonstrate the reduced memory usage, we present the peak memory during one step of decoding (after pre-fill has finished) in Table 4. In all cases, we observe a significant reduction in memory usage. This reduction is especially important on consumer hardware, such as on an nVidia RTX 3090 GPU where using QuaRot allows inference at 2048 sequence length which would be impossible otherwise.
Batch Size Speedup
1 1.97x
4 2.06x
16 2.11x
64 2.16x
Table 3: Time-to-first-token (prefill) speedup of each transformation block of Llama 2-7B model in QuaRot (over the FP16 model) on NVIDIA RTX 3090 GPU. We use 2048 sequence lengths with different batch sizes.
Batch Sequence Baseline QuaRot Saving
Size Length (GB) (GB) Factor
1 256 12.75 3.66 3.48x
512 12.88 3.69 3.49x
1024 13.13 3.76 3.49x
2048 13.64 3.89 3.50x
4096 14.72 4.22 3.49x
16 256 14.66 4.16 3.52x
512 16.69 4.69 3.56x
1024 20.76 5.76 3.60x
2048 28.88 7.88 3.66x
4096 45.20 12.19 3.71x
Table 4: Peak Memory usage (in GB) for decoding a single token using Llama 2-7B model with KV caches of different lengths and with different batch size.
@@ 5.3 Ablation Studies
To evaluate different aspects of QuaRot, we investigate the role of applying Hadamard transformation on the Weight-only Quantization schemes as well as using Random Orthogonal Matrices instead of Hadamard matrices. We also evaluate the accuracy of our quantized models when we apply FP16 Hadamard Transformation. At the quantization level, we evaluate the use of Round-to-Nearest Weight Quantization and Group-wise Quantization (with different group sizes). Finally, we evaluate using different precision for KV cache Quantization.
Weight-only Quantization.
@@
Random Orthogonal Matrices.
@@
FP16 Hadamard Transformation.
Table 7: Ablation on the precision of online Hadamard transformations for QuaRot. We use WikiText-2 perplexity as well as zero-shot tasks, explained in 5.3.
Precision
7B Baseline - 5.47 79.11 69.06 75.99 74.58 46.25 73.90 69.82
QuaRot FP32 6.10 76.77 63.77 72.16 69.87 40.87 70.39 65.64
FP16 6.08 76.99 66.46 72.59 69.07 41.21 70.59 66.21
13B Baseline - 4.88 80.47 72.22 79.39 77.48 49.23 76.75 72.59
QuaRot FP32 5.40 78.89 70.24 76.37 72.98 46.59 73.67 69.79
FP16 5.40 77.69 70.09 75.75 73.95 47.61 73.22 69.72
Round-to-Nearest Weight Quantization.
GPTQ is our default choice for weight quantization in QuaRot. Here, we study the role of quantizing the weights using Round-to-Nearest (RTN). Table 8 shows that applying GPTQ and RTN during the weight quantization maintains 99% and 98% respectively of the FP16 model accuracy in both 6 and 8 bits. In 4 bits, the gap between QuaRot-RTN and QuaRot-GPTQ decreases when the model size is increased (2.27 on Llama 2-7B and 0.34 on Llama 2-70B ) showing that GPTQ is a better option in smaller models.
Table 8: WikiText-2 Perplexity and zero-shot accuracy of QuaRot on the Llama- 2 family using 4, 6 and 8-bits with GPTQ and RTN weight quantization and RTN activation quantization. For zero-shot tasks, we use PIQA (PQ), WinoGrande (WG), HellaSwag (HS), Arc-Easy (A-e), Arc-Challenge (A-c), and LAMBADA (LA).
@@
Group-wise Quantization.
KV Cache Quantization.
We keep the rest of the model (including weights and activations) in high precision and apply our group-wise asymmetric quantization (with group-size 128) with various precision to keys and values. Table 10 shows the results of using various precision during KV cache quantization. The results show a negligible (at most 0.21) perplexity degradation up to 3-bit KV cache (0.07 for Llama 2-70B model). In addition, by comparing the 3 and 4-bit quantization, we can see that compared to the values, keys are more sensitive to quantization as keeping the keys in 4-bits and values in 3-bits has 0.03 perplexity loss (0.18 for 3-bit keys and 4-bit values) on the Llama 2-7B model. This matches the previous study on KV cache quantization (Hooper et al., 2024; Liu et al., 2024). The results show that using 3-bit KV-caches results in a better accuracy (5.68 on Llama 2-7B model) compared to keeping the keys in 4-bits and quantizing the values using 2-bits (with 5.75 perplexity on Llama 2-7B model).
Table 9: WikiText-2 perplexity of 4-bit QuaRot with various group-sizes on Llama- 2 models. We use GPTQ during the weight quantization. In all cases, we keep the KV cache group-size to 128 (which is the same as the head dimension).
Method Llama- 2
7B 13B 70B
Baseline 5.47 4.88 3.32
QuaRot 6.10 5.40 3.79
QuaRot-256G 5.98 5.28 3.63
QuaRot-128G 5.93 5.26 3.61
QuaRot-64G 5.88 5.25 3.58
K bits V bits Llama- 2
7B 13B 70B
16 16 5.47 4.88 3.32
4 4 5.51 4.91 3.33
4 3 5.54 4.93 3.35
4 2 5.75 5.09 3.43
3 4 5.65 5.01 3.38
3 3 5.68 5.02 3.39
3 2 5.93 5.21 3.48
2 4 8.06 6.42 3.89
2 3 8.18 6.50 3.92
2 2 9.23 7.07 4.13
6 Conclusion
We introduce QuaRot: a method which uses Hadamard matrices to eliminate outliers in the activations and KV cache of pre-trained LLMs, enabling end-to-end 4-bit quantization for the first time (to the best of our knowledge). Quantizing Llama 2-70B to 4 bits with QuaRot maintains 99% of the downstream task performance of the FP16 baseline, with a 2.16x speedup on RTX 3090 GPUs during the prefill stage (and up to 3.39x memory saving during the decoding stage). Quantizing all Llama- 2 models to 6- and 8-bits is lossless.
Appendix A Appendix
A.1 Clipping Ratio
1.0 0.95 0.9 0.85
Input Quantization 5.938 5.910 5.828 5.850
KV Cache Quantization 5.513 5.510 5.517 5.532
A.2 Performance Analysis: Detailed Results

### v2

@@ 5 Experimental Validation
We implement QuaRot using Hugging Face (Wolf et al., 2019) on top of the PyTorch framework (Paszke et al., 2019). To quantize the inputs, we use per-token symmetric quantization (a single scale for every row) with a constant clipping ratio of 0.9 in all our experiments. We quantize the KV caches using asymmetric quantization with a group size 128 with a constant clipping ratio of 0.95. For weight quantization, we use round-to-nearest (RTN) and GPTQ (Frantar et al., 2022) with per-column (also known as per-channel) symmetric quantization, where we extract the clipping ratio using a linear search over the squared error. We use 128 samples from WikiText-2 (Merity et al., 2016) training set with 2048 sequence length as the calibration set during GPTQ quantization. On a single NVIDIA A100 GPU, modifying Llama 2-70B with QuaRot takes 5 minutes and quantizing the model with GPTQ takes a further 2 hours. We present Llama- 3 results in Appendix A.8.
@@ 5.1 Accuracy Results
First, we evaluate the accuracy of QuaRot on the language generation task. Table 1 shows the perplexity of Llama- 2 models on WikiText-2 when we quantize the weights using GPTQ. We compare against 4-bit SmoothQuant (Xiao et al., 2023) and OmniQuant (Shao et al., 2023). We also include the QUIK (Ashkboos et al., 2023) results when they keep all the layers (including down-projection) in 4 bits. QuaRot outperforms all previous work with at most 0.63 perplexity loss (0.47 on Llama 2-70B model) without any re-training (as in OmniQuant) nor higher precision outlier features and asymmetric quantization (as in QUIK). We also apply group-wise quantization to compare against Atom (Zhao et al., 2023) on the same number of groups for weight and activations. In this setting, QuaRot doesn’t need to keep any higher precision features and related operations (like re-ordering). QuaRot outperforms Atom with 0.1 perplexity points in the 7B model. On the 13B model, we get the same perplexity number as Atom.
Table 1: WikiText-2 perplexity results on 4-bit quantization of Llama- 2 models with 2048 sequence length. We extract the results for SmoothQuant and OmniQuant results of (Shao et al., 2023). 128G shows the group-wise quantization with group size 128.Here, we quantize all weights, activations, and caches in 4-bits in QuaRot.
@@
Next, we focus on evaluating QuaRot on six important zero-shot tasks: PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), LAMBADA (OpenAI) (Radford et al., 2019), and Arc (Easy and Challenge) (Clark et al., 2018). We use the LM Evaluation Harness (Gao et al., 2021) with default parameters for our experiments. Table 2 shows the accuracy of our scheme on the above tasks as well as the average score. On Llama- 2 family, QuaRot preserves the accuracy with at most 4.18% average score loss (1.09% for 70B model).
Table 2: Zero-shot accuracy of Llama- 2 models with 4-bit (A4W4KV4) QuaRot on PIQA (PQ), WinoGrande (WG), HellaSwag (HS), Arc-Easy (A-e), Arc-Challenge (A-c), and LAMBADA (LA).
@@ 5.2 Performance Analysis
We implement QuaRot using CUDA/12.1 on top of PyTorch and use CUTLASS for performing INT-4 matrix multiplication on TensorCore (where the results will be saved in an INT32 accumulator). In this section, we evaluate the performance of our kernels for both prefill and decoding steps on NVIDIA RTX 3090 GPU. We provide all our experiments on a single transformer block as the whole model does not fit on our GPU cluster for large batch sizes. We provide more performance analysis of our kernels (as well as complete results) in Appendix A.10.
Figure 4: Performance of the QuaRot kernel on a single transformer block of Llama- 2 models using NVIDIA RTX 3090 GPU. Left: For the speedup results, we evaluate using sequence length 2048 with different batch sizes. Right: Peak memory saving during decoding of 50 tokens with different prefill sequence lengths using batch size 16.
Prefill Stage Performance Increases.
For the compute-bound prefill stage, we present the speedups of using QuaRot on 2048 sequence length with different batch sizes in Figure 4 Left. On Llama 2-7B model, we get 1.97x-2.16x speedup over the FP16 implementation using our QuaRot kernel. The speedup increases with batch sizes as the computation will become a bottleneck in larger batch sizes. on Llama 2-70B model, we get up to 3.33x speedup. Note that our performance results could be improved by optimizing our kernels (e.g., fusing the quantization operations into the MatMul).
Decoding Stages Memory Saving.
Finally, we evaluate the memory improvement which is the main bottleneck of the decoding stage. Figure 4 Right shows the peak memory saving on Llama- 2 models. We provide results for Llama 2-7B and Llama 2-70B models. In both models, we get at least 3.63x peak memory saving compared to FP16 case during the decoding stage. Note that the KV cache is larger in Llama 2-7B model as the Llama 2-70B uses grouped-query attention (Ainslie et al., 2023). In the Llama 2-7B model, the memory saving increases with the sequence length, resulting in up to 3.75x memory saving. on Llama 2-70B model, we get 3.89x savings in almost all the cases. We expect these values to be larger for the whole model (instead of just the single layer here) since as the number of layers increases the effect of constant size objects in memory becomes much less significant.
@@ 5.3 Ablation Studies
To evaluate different aspects of QuaRot, we evaluate the use of Round-to-Nearest Weight Quantization, Group-wise Quantization (with different group sizes), and KV cache Quantization with different bit-width combinations (Appendix A.3). In addition, we investigate the role of applying Hadamard transformation on the Weight-only Quantization schemes (Appendix A.4) as well as using Random Orthogonal Matrices (Appendix A.5) instead of Hadamard matrices. Finally, we evaluate the accuracy of our quantized models when we apply FP16 Hadamard Transformation (Appendix A.7).
Round-to-Nearest Weight Quantization.
GPTQ is our default choice for weight quantization in QuaRot. Here, we study the role of quantizing the weights using Round-to-Nearest (RTN). Table 3 shows that applying RTN weight quantization fully maintains the FP16 model accuracy in 8 bits. We note that RTN does not need any calibration set or hyper-parameter during the quantization. Comparing Table 3 and 2, we conclude that in 4 bits, the gap between QuaRot-RTN and QuaRot-GPTQ decreases when the model size is increased (2.27 on Llama 2-7B and 0.34 on Llama 2-70B ) showing that GPTQ is a better option in smaller models. For more detailed results see Appendix A.6.
Table 3: WikiText-2 Perplexity and zero-shot accuracy of QuaRot on the Llama- 2 family using 4- and 8-bits with Round-to-Nearest (RTN) weights and activation quantization. For zero-shot tasks, we use PIQA (PQ), WinoGrande (WG), HellaSwag (HS), Arc-Easy (A-e), Arc-Challenge (A-c), and LAMBADA (LA). We quantize all weights, activations, and caches.
7B Baseline FP16 5.47 79.11 69.06 75.99 74.58 46.25 73.90 69.82
QuaRot-RTN INT4 8.37 72.09 60.69 65.40 58.88 35.24 57.27 58.26
INT8 5.50 78.94 68.67 75.80 74.79 45.39 74.33 69.65
70B Baseline FP16 3.32 82.70 77.98 83.84 80.98 57.34 79.58 77.07
QuaRot-RTN INT4 4.14 80.69 75.14 79.63 77.57 51.71 77.02 73.63
INT8 3.33 82.97 77.98 83.67 80.77 58.11 79.53 77.17
Group-wise Quantization.
Table 4: WikiText-2 perplexity of 4-bit QuaRot with various group-sizes on Llama- 2 models. We use GPTQ during the weight quantization. In all cases, we keep the KV cache group-size to 128 (same as the head dimension). 128G shows the group-wise quantization with 128 group size.
Method Llama- 2
7B 13B 70B
Baseline 5.47 4.88 3.32
QuaRot 6.10 5.40 3.79
QuaRot-256G 5.98 5.28 3.63
QuaRot-128G 5.93 5.26 3.61
QuaRot-64G 5.88 5.25 3.58
6 Conclusion
We introduce QuaRot: a method which uses Hadamard matrices to eliminate outliers in the activations and KV cache of pre-trained LLMs, enabling end-to-end 4-bit quantization for the first time (to the best of our knowledge). Quantizing Llama 2-70B to 4 bits with QuaRot maintains 99% of the downstream task performance of the FP16 baseline, with a 2.16$\times$ speedup on RTX 3090 GPUs during the prefill stage (and up to 3.39$\times$ memory saving during the decoding stage). Quantizing all Llama- 2 models to 6 and 8 bits is lossless.
Appendix A Appendix
A.1 QuaRot on the Attention Module
Figure 5 shows the original attention module in large language models with RoPE. The input of the attention module is already rotated using the randomized Hadamard matrix ${\mathbf{Q}}$ (see Section 4) and in the first step, we fuse the inverse of such matrices into the input linear layers of the attention. In the next step, we fuse the exact Hadamard matrices on each block of the columns (proportional to each head) on the V_projection layer to make sure that the Values will be rotated at the output of that layer. In the next step, we apply exact Hadamard transformations on the Keys and Queries and quantize the KV after RoPE operation (note that the Keys and Queries Hadmard transformations will be canceled during the attention operation). Finally, we apply another Hadamard transformation between heads before Out_projection layer and fuse the inverse into the weights. Figure 6 shows the result of applying QuaRot on the attention module.
[…]
A.2 Clipping Ratio Ablation
1.0 0.95 0.9 0.85
Input Quantization 5.938 5.910 5.828 5.850
KV Cache Quantization 5.513 5.510 5.517 5.532
A.3 KV Cache Quantization Ablation
We keep the rest of the model (including weights and activations) in high precision and apply our group-wise asymmetric quantization (with group-size 128) with various precision to keys and values. Table 6 shows the results of using various precision during KV cache quantization. The results show a negligible (at most 0.21) perplexity degradation up to 3-bit KV cache (0.07 for Llama 2-70B model). In addition, by comparing the 3 and 4-bit quantization, we can see that compared to the values, keys are more sensitive to quantization as keeping the keys in 4-bits and values in 3-bits has 0.03 perplexity loss (0.18 for 3-bit keys and 4-bit values) on the Llama 2-7B model. This matches the previous study on KV cache quantization [Hooper et al., 2024, Liu et al., 2024]. The results show that using 3-bit KV-caches results in a better accuracy (5.68 on Llama 2-7B model) compared to keeping the keys in 4-bits and quantizing the values using 2-bits (with 5.75 perplexity on Llama 2-7B model).
K bits V bits Llama- 2
7B 13B 70B
16 16 5.47 4.88 3.32
4 4 5.51 4.91 3.33
4 3 5.54 4.93 3.35
4 2 5.75 5.09 3.43
3 4 5.65 5.01 3.38
3 3 5.68 5.02 3.39
3 2 5.93 5.21 3.48
2 4 8.06 6.42 3.89
2 3 8.18 6.50 3.92
2 2 9.23 7.07 4.13
A.4 Weight-only Quantization Ablation
@@
A.5 Random Orthogonal Matrices Ablation
@@
A.6 Round-to-Nearest Weight Quantization: Detailed Results
Table 9 shows the detailed results of QuaRot with GPTQ and round-to-nearest (RTN) weight quantization for both 6 and 8 bits on various tasks for Llama- 2 models.
Table 9: WikiText-2 Perplexity and zero-shot accuracy of QuaRot on the Llama- 2 family using 4, 6 and 8-bits with GPTQ and RTN weight quantization and RTN activation quantization. For zero-shot tasks, we use PIQA (PQ), WinoGrande (WG), HellaSwag (HS), Arc-Easy (A-e), Arc-Challenge (A-c), and LAMBADA (LA).The Precision column shows the bitwidth for all inputs, weights, and KV-caches.
@@ A.6 Round-to-Nearest Weight Quantization: Detailed Results
A.7 FP16 Hadamard Transformation Ablation
Table 10: Ablation on the precision of online Hadamard transformations for QuaRot. We use WikiText-2 perplexity as well as zero-shot tasks, explained in Section 5.3.
Precision
7B Baseline - 5.47 79.11 69.06 75.99 74.58 46.25 73.90 69.82
QuaRot FP32 6.10 76.77 63.77 72.16 69.87 40.87 70.39 65.64
FP16 6.08 76.99 66.46 72.59 69.07 41.21 70.59 66.21
13B Baseline - 4.88 80.47 72.22 79.39 77.48 49.23 76.75 72.59
QuaRot FP32 5.40 78.89 70.24 76.37 72.98 46.59 73.67 69.79
FP16 5.40 77.69 70.09 75.75 73.95 47.61 73.22 69.72
A.8 Llama- 3 Results
In this section, we show the accuracy of applying QuaRot for quantizing the Llama 3-8B and Llama 3-70B models. Table 11 shows the WikiText-2 perplexity of quantizing the Llama- 3 models with QuaRot using 4-bit quantization. Compared to Table 1, we conclude that Llama- 3 is more sensitive to quantization as we can see a higher gap between the quantized and FP16 models. Table 12 shows the accuracy results of those models on zero-shot tasks.
Table 11: WikiText-2 perplexity results on 4-bit quantization of Llama- 3 models with 2048 sequence length. 128G shows the group-wise quantization with group size 128.
Method Weight #Outlier Llama- 3
Quantization Features 8B 70B
Baseline - - 6.14 2.86
QuaRot GPTQ 0 8.16 6.66
QuaRot-128G GPTQ-128G 0 7.36 5.51
Table 12: Zero-shot accuracy of Llama- 3 models with 4-bit QuaRot on PIQA (PQ), WinoGrande (WG), HellaSwag (HS), Arc-Easy (A-e), Arc-Challenge (A-c), and LAMBADA (LA).
Model Method PQ WG HS A-e A-c LA Avg.
Llama 3-8B FP16 80.74 72.77 79.06 77.82 53.33 75.63 73.22
QuaRot 75.14 65.82 72.94 68.01 43.34 65.81 65.18
Llama 3-70B FP16 84.66 80.51 84.89 85.86 64.25 79.47 79.94
QuaRot 78.07 69.30 77.33 73.44 47.53 69.57 69.21
A.9 Phi-3-mini-4k-instruct Results
In this section, we show the accuracy of applying QuaRot for quantizing the Phi-3-mini-4k-instruct model [Abdin et al., 2024]. Table 13 shows the accuracy results of the model in terms of perplexity and on zero-shot tasks.
Table 13: WikiText-2 Perplexity and zero-shot accuracy of QuaRot on the Phi-3-mini-4k-instruct model (revision = ff07dc01) using 4, 6 and 8-bits with GPTQ and RTN weight quantization and RTN activation quantization. For zero-shot tasks, we use PIQA (PQ), WinoGrande (WG), HellaSwag (HS), Arc-Easy (A-e), Arc-Challenge (A-c), and LAMBADA (LA).
Phi-3-mini Baseline FP16 6.35 80.47 73.72 78.45 80.13 57.51 68.37 73.11
QuaRot-RTN INT4 11.69 68.39 58.64 60.60 65.87 39.25 43.99 56.12
INT6 6.78 79.54 73.01 77.46 79.21 55.12 67.53 71.98
INT8 6.58 79.71 74.11 78.63 80.47 56.66 68.56 73.02
QuaRot-GPTQ INT4 7.85 75.35 67.88 72.95 72.98 48.12 60.78 66.34
INT6 6.63 79.54 72.69 78.50 79.42 56.74 68.85 72.67
INT8 6.58 80.25 74.19 78.54 80.35 57.08 68.64 73.18
A.10 Performance Analysis
We implement the attention mechanism using three routines: 1) Init: During the prefill stage, this routine initializes the cache from all the key and value vectors in the prefill. The attention output during prefill is computed directly using Flash Attention [Dao et al., 2022] since we already have access to dequantized keys and values. 2) Append: During decoding, this routine is called first to quantize the current keys and values and append them to the cache. 3) Decode: Finally, this routine is called during decoding with the current query vector. The routine computes the attention output using a quantized implementation of flash attention which can load the quantized cache and compute the final value vector.
4-bit Linear and Attention Layers.
We benchmark our 4-bit linear layer which involves 4-bit matrix multiplication. For a given input of FP16, the layer optionally computes the Hadamard operation, then calls the quantization kernel to quantize and save the input in a sub-byte format. In the next step, the quantized weights and input are passed to the CUTLASS 4-bit GEMM kernel. Finally, the output is dequantized and cast back to FP16. Figure 7 shows the speedup of our 4-bit layer for different layer sizes where the layer sizes match the FFN linear layer sizes in Llama- 2 models. Our 4-bit linear layer gets 3.2x speedup relative to FP16 in the Llama 2-7B model, and 4.3x on the Llama 2-70B model. These numbers are for a batch size of 1, we find that scaling is approximately linear with batch size: more results in Table 14. We include the runtime with and without Hadamard operations, as ${\mathbf{W}}_{\textrm{up}}$ and ${\mathbf{W}}_{\textrm{gate}}$ do not require Hadamard transforms, whilst ${\mathbf{W}}_{\textrm{down}}$ does. We see that the Hadamard transform adds very little overhead to the forward pass at most 7% overhead.
Figure 7: Performance of 16-bit and 4-bit linear layer for 2048 sequence lengths with and without online Hadamard transformation on a NVIDIA RTX 3090 GPU, averaged over 1000 runs. The matrix sizes correspond to the linear layer sizes in Llama- 2 FFN blocks (i.e. ${\mathbf{W}}_{\textrm{down}}$). Here the batch size is 1, but the performance ratio holds for larger batches (see Table 14).